OpenStudy (anonymous):

how do i solve 3x over 5 - x = x over 10 - 5 over 2

OpenStudy (anonymous):

the minus x is not under the 3x over 5

OpenStudy (precal):

type it in using the equation editor

OpenStudy (anonymous):

idk how to do that... it's just 3x/5 then - x = x/10 - 5/2

OpenStudy (anonymous):

\[\frac{3x}{5} - x = \frac{x}{10} - \frac{5}{2}\]???

OpenStudy (anonymous):

yes like that !

OpenStudy (anonymous):

OK, so: 3x/5 - x = x/10 - 5/2 let's just multiply both sides by 10: 6x - 10x = x - 25 Then combine like terms: -4x = x - 25 -5x = -25 x = 5

OpenStudy (anonymous):

okay thank you ! can i just ask , that , for problems like this , you multiply by the least common denominator , right ? or ?

OpenStudy (anonymous):

Sure. That's just to make the math easier to do, though. You could also just solve it directly: 3x/5 - x = x/10 - 5/2 (3/5)x - (5/5)x = x/10 - 5/2 -(2/5)x = x/10 - 5/2 -(4/10)x = (1/10)x - 5/2 -(5/10)x = -5/2 -x/2 = -5/2 x = 5

OpenStudy (anonymous):

okay thanks !